Light emitting diodes (LED) light bulbs have become required in recent years, but do they make financial sense? Suppose a typical 60-watt incandescent light bulb costs $.42 and lasts 1,000 hours. A 15-watt LED, which provides the same light, costs $3.25 and lasts for 12,000 hours. A kilowatt-hour of electricity costs $118. A kilowatt-hour is 1,000 watts for 1 hour. If you require a return of 11 percent and use a light fixture 500 hours per year, what is the equivalent annual cost of each light bulb? (A negative answer should be ind icated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) 60-watt incandescent light bult LED light bulb

I need solution
Answer #1
A kilowatt hour is 1000 watts for 1 hour
60 watt bulb burning for 500 hours per year uses 30000 watt hours or 30 kilowatt hours
Kilowatt electricity cost is $0.118 the cost per year would be (30*0.118) 3.54
NPV of 60 watt is -0.42 - 3.54*PVIFA(11%,2)
-6.48402
PVIFA = 1.713
EAC = -6.48402/1.713 -3.79
EAC for LED Light Bulb
15 watt bulb burning for 500 hours per uses (15*500) = 7500 watt or 7.5 kilowatt hours
Kilowatt electricity cost is $ 0.118, cost per year would be (7.5*0.118) 0.885
12000/500 24
NPV of 15 watt is -3.25 - 0.885*PVIFA(11%,24)
NPV -10.638
EAC = 10.638/8.348 -1.27
PVIFA = 8.348